OpenStudy (alex6799):

PLEASE HELP!!! Find a quadratic model for the set of values: (-2, -20), (0, -4), (4, -20). Show your work.

OpenStudy (solomonzelman):

\(\large\color{#000000}{\displaystyle y= ax^2+bx+c }\) is what the quadratic equation looks like.

OpenStudy (solomonzelman):

I got disconnected, apologize.

OpenStudy (alex6799):

ok and then what? i have no idea on how to do this :(

OpenStudy (solomonzelman):

You are given 3 points so plug them in (individually) to solve for a, b, & c.

OpenStudy (alex6799):

oh no worries its fine

OpenStudy (solomonzelman):

plug in the points; For (-2,-20), \(\large\color{#000000}{\displaystyle -20= a(-2)^2+b(-2)+c }\) For (0,-4), \(\large\color{#000000}{\displaystyle -4= a(0)^2+b(0)+c }\) For (4,-20), \(\large\color{#000000}{\displaystyle -20= a(4)^2+b(4)+c }\)

OpenStudy (solomonzelman):

Simplify and solve the system...

OpenStudy (solomonzelman):

(note: it is very easy to obtain the c from the 2nd equation)

OpenStudy (alex6799):

i dont understand

OpenStudy (solomonzelman):

each of those points (-2, -20), (0, -4), (4, -20) is on the parabola, therefore, you can plug them into the parabola to solve for a, b and c. I plugged everything for you you just need to simplify and solve.

OpenStudy (solomonzelman):

let's go with the first equation: \(\large\color{#000000}{\displaystyle -20= a(-2)^2+b(-2)+c }\) can you simplify this one?

OpenStudy (alex6799):

ok i get that but how would i simplify them?

OpenStudy (solomonzelman):

(-2)² = ?

OpenStudy (alex6799):

4?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{#000000}{\displaystyle -20= a(-2)^2+b(-2)+c }\) \(\large\color{#000000}{\displaystyle -20= 4a-2b+c }\)

OpenStudy (solomonzelman):

\(\large\color{#000000}{\displaystyle -20= 4a-2b+c }\) is your first equation.. ok?

OpenStudy (alex6799):

ok

OpenStudy (solomonzelman):

Now simplify the second equation

OpenStudy (alex6799):

\[-4=a+b+c\] ?

OpenStudy (alex6799):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{#000000}{\displaystyle -4= a(0)^2+b(0)+c }\) what is 0²?

OpenStudy (alex6799):

0 right?

OpenStudy (solomonzelman):

yes, and what is a times 0?

OpenStudy (alex6799):

0

OpenStudy (solomonzelman):

what is b•0 ?

OpenStudy (alex6799):

0?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So what does the second equatrion simplify to?

OpenStudy (alex6799):

-4=a+c?

OpenStudy (solomonzelman):

almost right

OpenStudy (solomonzelman):

we said that 0²=0, and a•0=0, so what happens to the "a(0)²" component?

OpenStudy (alex6799):

it goes away?

OpenStudy (solomonzelman):

yes, so what is your second equation?

OpenStudy (alex6799):

umm... -4=c?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{#000000}{\displaystyle -20= 4a-2b+c }\) \(\large\color{#000000}{\displaystyle -4= c }\) now lets do the 3rd equation.

OpenStudy (solomonzelman):

\(\large\color{#000000}{\displaystyle -20= a(4)^2+b(4)+c }\)

OpenStudy (solomonzelman):

4² = ?

OpenStudy (alex6799):

8?

OpenStudy (alex6799):

wait no 16

OpenStudy (solomonzelman):

yes 4² = 4•4 = 16

OpenStudy (solomonzelman):

so your third equation, when simplified, would be?

OpenStudy (alex6799):

\[-20=16a+4b+c\] ?

OpenStudy (solomonzelman):

fabulous!

OpenStudy (alex6799):

^-^

OpenStudy (solomonzelman):

So we have, \(\large\color{#000000}{\displaystyle -20= 4a-2b+c }\) \(\large\color{#000000}{\displaystyle -4= c }\) \(\large\color{#000000}{\displaystyle -20=16a+4b+c }\)

OpenStudy (solomonzelman):

Note (again), that in 2nd equation you are given the value of c explicitely: c=-4 And you can use that to solve the 1st and 2rd equations for "a" and "b".

OpenStudy (solomonzelman):

plug in c=-4, into equation 1 and into equation 3.

OpenStudy (alex6799):

so... \[-20=4a-2b+-4\] ?

OpenStudy (solomonzelman):

yes, the first equation... and that would be better written as, \(\large\color{#000000}{\displaystyle -20=4a-2b-4 }\)

OpenStudy (solomonzelman):

Now, plug in c=-4, into the 3rd equation, please..

OpenStudy (alex6799):

\[-20=16a+4b+-4\] ?

OpenStudy (solomonzelman):

YES! and that would be better written as?

OpenStudy (alex6799):

@SolomonZelman ?